If f and g are differentiable functions for all real values of x such that f(1) = 4, g(1) = 3, f '(3) = −5, f '(1) = −4, g '(1) = −3, g '(3) = 2, then find h '(1) if h(x) = f(x) g(x).
−9
−24
0
24

Answer by ikleyn(53763)

.
If f and g are differentiable functions for all real values of x such that f(1) = 4, g(1) = 3, f '(3) = −5, f '(1) = −4, g '(1) = −3, g '(3) = 2, then find h '(1) if h(x) = f(x) g(x).
−9
−24
0
24
--------------------------------------------------------------------
h'(1) = f'(1)*g(1) + f(1)*g'(1) = (-4)*3 + 4*(-3) = -24.